### v1

# How segmented is my network?

Abstract
Network segmentation is a popular security practice for limiting lateral movement, yet practitioners lack a metric to measure how segmented a network actually is. We model a network as a graph and study segmentedness as a property captured by the global edge density that can be estimated from sampled node pairs. Then, we derive an estimator and evaluate its uncertainty using confidence intervals. For a 95% confidence interval with a margin-of-error of $\pm 0.1$, we show that a minimum of $M=97$ sampled node pairs is sufficient. This result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random. We validate the estimator through Monte Carlo simulations on Erdős–Rényi and stochastic block models, demonstrating accurate estimation and well-behaved coverage. Finally, we discuss applications of the estimator, such as, baseline tracking, zero trust assessment, and merger integration.
I Introduction
Network segmentation is a widely recommended security practice for limiting lateral movement and reducing the impact of breaches. Modern security frameworks routinely advise practitioners to divide networks into smaller, isolated zones and tightly control communication between them. Despite this emphasis, security teams still lack a simple way to assess how segmented a network is. In practice, segmentation is usually assessed qualitatively using architecture diagrams, policy reviews, or expert judgment.
The lack of a quantitative measure has concrete consequences. Organizations cannot objectively compare segmentation across business units, validate that architecture changes improved isolation, or determine whether network complexity is justified by security benefits. Without measurement, segmentation remains a qualitative aspiration rather than a managed property. Further, practitioners are left without objective evidence to support segmentation decisions or investments.
In this paper, we model a network as a graph and study segmentedness as a measurable property of that graph. We show that a simple statistic derived from the network’s global edge density captures meaningful information about network flatness. We provide a principled method to estimate this quantity from observed data and attach confidence intervals to the estimate. Our approach is intentionally lightweight and does not require detailed knowledge of application semantics or traffic intent. Through analysis and simulation, we demonstrate that the estimator is accurate and well-behaved across a range of network models. We argue that this provides security practitioners with a practical, interpretable metric for reasoning about segmentation.
The rest of the paper is organized as follows. Section II reviews prior work on network segmentation and graph-based metrics. Section III introduces our network model and formally defines segmentedness in terms of global edge density. Section IV presents our estimator and derives confidence intervals for segmentedness from observed network data. Section V provides example sample size calculations for practical deployment. Section VI evaluates the estimator using synthetic networks, including Erdős–Rényi and stochastic block models. Section VII discusses some operational use cases using the segmentedness metric. Section VIII discusses practical considerations for sampling in operational environments. Section IX makes explicit the limitations of our approach. Section X concludes with a summary of findings and directions for future work.
II Related Work
We review prior work on network segmentation and graph-based measures potentially related to segmentedness. None of the works reviewed below set out to create a segmentedness (or flatness) measure for security practitioners. As such, our review should be seen as a summary of prior research focus.
II-A Government and Industry Frameworks
NIST SP 800-207 [18] defines network segmentation and micro-segmentation as core design principles of zero trust architectures, emphasizing that access between resources should be tightly controlled and continuously verified. However, the framework describes segmentation in architectural and qualitative terms, such as the presence of trust boundaries and enforcement points. As a result, segmentation assessment relies on expert judgment, not on a quantitative measure of the degree of segmentedness or network flatness.
Similarly, CISA Zero Trust Maturity Model [6] defines maturity across a four-stage ordinal scale from “traditional” to “optimal” implementations. While this model provides a strategic roadmap, it relies on qualitative self-assessment and high-level checklists to determine a network’s maturity level. Consequently, an organization’s placement on this scale remains a subjective determination rather than a reflection of a verifiable metric.
NIST CSF 2.0 [15] treats network segmentation as an important security control within its broader guidance for managing cybersecurity risk. The framework evaluates segmentation through qualitative categories and maturity-based profiles, where organizations assess whether network security policies are in place. CSF 2.0 also relies on expert self-assessment and descriptive tiers, making it difficult to quantitatively compare segmentation strength over time or across organizations.
Taken together, these government and industry frameworks emphasize the importance of segmentation but assess it through qualitative judgment, highlighting the absence of a quantitative, repeatable scalar measure for evaluating how segmented a network is or how that segmentedness changes over time.
II-B Network Segmentation Research
[21] evaluates network segmentation by measuring how different segmentation designs affect both security and the ability of a network to carry out its intended operational tasks. The paper introduces a unified quantitative metric that combines a security score, representing how often devices remain usable during an attack, with a delay score, representing how much normal network operations are slowed by the attack. This metric therefore, measures the practical effectiveness of segmentation in maintaining secure and timely network operation. It does not directly measure how segmented or flat the network is.
[19] introduces a quantitative Segmentation Index (SI) that is used to evaluate and select a good way of dividing a network into segments. The index measures how well machines within the same segment belong together and how clearly different segments are separated by firewall rules. However, SI does not (and was not designed to) indicate how flat or segmented the network is overall.
[17] and [3] both evaluate network micro-segmentation through its security effects rather than by defining a single direct measure of segmentedness. [17] compares different segmentation setups in smart home IoT networks and shows that stronger segmentation leads to reduced attack surface, lower vulnerability exposure, and fewer compromised devices. Similarly, [3] assesses segmentation by measuring how it reduces network exposure and attacker capabilities using multiple graph-based metrics such as connectivity, reachability, centrality, and attack-path characteristics. In both works, segmentedness is inferred from improved security. Segmentedness is not captured by a single, scalar metric.
[13] gives a precise mathematical definition of network segmentation based on how similar or different machines’ access-control rules are, enabling automated construction and validation of segmentation designs. Under this framework, a network is treated as either correctly segmented or not. While this binary view is useful for formal reasoning, it does not reflect real networks, where segmentation is often uneven and incremental rather than all-or-nothing.
[20] introduces LMSk as a family of metrics parameterized by the maximum number of attacker steps $k$, enabling analysis of how attacker effectiveness increases as deeper lateral movement is permitted. For a fixed $k$, LMSk is computed by estimating, for every pair of distinct machines, the likelihood that an attacker who compromises one can reach the other by chaining together up to $k$ plausible movement steps, with each step weighted by the ease with which the underlying connections can be abused, and then averaging these likelihoods across all pairs. Lower values of LMSk correspond to networks in which compromise is difficult to propagate under the assumed attacker model, while higher values indicate environments in which lateral movement is broadly feasible once an initial foothold is obtained. Although LMSk was not proposed as a direct measure of segmentation, it provides useful insight into the security consequences of network structure under varying attacker capabilities. At the same time, the dependence on an explicit attacker model, the requirement to choose $k$, and the absence of a single scalar summary limit its suitability as a standalone measure of segmentedness.
In [5], the authors experimentally demonstrate the security benefits of segmentation by comparing a flat network architecture against a segmented design during a simulated attack. Their results quantify the reduction in successful lateral movement when firewalls are introduced between zones. However, the methodology treats segmentation as a binary condition. Thus, the approach validates the utility of segmentation, but does not provide a mechanism to measure the relative degree of isolation in more complex, partially segmented environments.
[14] evaluates micro-segmentation by measuring its impact on network performance, showing that enabling segmentation adds only minimal delay and no packet loss. [11] discusses network segmentation as a security best practice and evaluates it qualitatively through design principles and case studies. [23] focuses on automatically creating segments in IIoT networks using traffic clustering and measures the quality of the clustering. [2] compares different segmentation and micro-segmentation deployment approaches using high-level operational factors such as security level, cost, and complexity. [12] evaluates role-based micro-segmentation in public clouds using metrics like role inference accuracy and policy violations, focusing on whether segmentation rules are applied correctly rather than on how segmented the network is. [7] provides a comprehensive taxonomy of segmentation concepts and practices across industry and research, documenting the diverse ways in which segmentation is defined and measured, and observes the lack of a generally accepted scalar metric for measuring the degree of segmentation or network flatness.
In summary, while recent literature has transitioned from qualitative best practices to quantitative evaluation, a significant gap remains. The most advanced current approaches, such as those in [3] and [20], provide measures of risk such as exposure, robustness, and lateral-movement susceptibility. However, these are fundamentally outcome-driven metrics that model the probability of specific attack behaviors within a given topology. To our knowledge, no prior work has proposed a scalar metric that quantifies “segmentedness” as an intrinsic topological property—independent of specific threat models or assumptions about how likely individual services are to be compromised. Such a metric would allow practitioners to answer the fundamental question, “How segmented is my network?” by providing a direct measure of network segmentedness (flatness) that can be tracked consistently across different designs and over time.
II-C Graph-based Measures
While much of the prior work on network segmentation treats segmentation as a binary architectural or policy correctness property, graph-based measures provide quantitative characterizations of network structure that are related to segmentation as understood in security practice. These measures capture properties such as structural cohesion and separability, which are relevant to understanding certain aspects of network organization and risk. They have been applied to graph analysis and represent natural candidates for quantifying network segmentation. However, as discussed below, thus far they have not been adapted to directly quantify the overall permissiveness of communication in a network.
Edge density is a simple graph-based measure [22] that quantifies how many connections exist in a graph relative to how many connections are possible. Given a graph with $n$ nodes, there are $\binom{n}{2}$ possible pairwise connections, and the edge density is defined as the fraction of those possible connections that are present. An edge density of one corresponds to a fully connected graph in which every node can communicate with every other node, while an edge density of zero corresponds to a graph with no allowed communication between nodes. Because it directly reflects the proportion of permitted communication relationships, edge density provides an intuitive measure of how flat (interconnected) a network is, independent of how those connections are arranged. Although prior work has not proposed edge density as a segmentedness metric, it provides a natural foundation for such a measure.
Modularity is another graph-based measure that evaluates how well a graph can be divided into groups, or communities, whose nodes are more densely connected to each other than to the rest of the graph [16]. Intuitively, modularity compares the number of connections within proposed groups to the number that would be expected if connections were placed at random, while preserving overall connectivity levels. High modularity values indicate that the graph has a clear community structure, with relatively strong separation between groups, whereas low modularity values suggest that connections are more evenly distributed and that the graph lacks meaningful subdivision. As such, modularity is useful for identifying and evaluating segmentation in terms of clustered structure, but it emphasizes the arrangement of connections. In contrast, edge density directly reflects how many communication relationships are permitted overall, independent of how those relationships are grouped.
The Fiedler value, also known as algebraic connectivity, is a numerical property of a graph that captures how strongly connected the graph is as a whole and how difficult it is to separate into disconnected components [9]. A value of zero indicates that the graph is disconnected, while larger values indicate increasing resistance to separation, with graphs containing bottlenecks or hub-and-spoke structures typically exhibiting low Fiedler values. As such, the Fiedler value is relevant to network segmentedness when segmentedness is interpreted in terms of structural separability or vulnerability to cuts. However, it is less suitable when segmentedness is interpreted as the overall permissiveness of communication, since graphs with the same number of allowed connections can have very different Fiedler values depending on how those connections are arranged.
Despite their relevance to network communications, these measures have not been turned into segmentedness metrics for security practitioners.
III Formal Model
We propose defining segmentedness as the overall permissiveness of communication policy, formalized through edge density. Accordingly, we model segmentedness using edge density, which directly captures the fraction of end-to-end communication relationships that are permitted and serves as the foundation for the formal model developed in the remainder of this section.
We acknowledge that the structural arrangement of permitted connections can influence security outcomes, for example, by creating chokepoints or enabling indirect communication paths. However, as documented in Section II, security practitioners currently lack a simple quantitative metric that summarizes how permissive a network’s communication policy is; our objective prioritizes a single, easily interpretable scalar over metrics that emphasize structure. This objective explicitly rules out structure-oriented measures such as modularity and the Fiedler value, which capture how connections are arranged rather than how much communication is permitted overall.
III-A Network Model
We model a network as an undirected simple graph $G=(V,E)$, where $V$ is the set of nodes representing end systems such as user devices or server workloads, and $E\subseteq\{\{u,v\}\mid u,v\in V,u\neq v\}$ is the set of edges. An edge $\{u,v\}\in E$ indicates that end-to-end communication between nodes $u$ and $v$ is permitted by network security policy, independent of the physical or logical path taken through the underlying network. A missing edge between two nodes means that direct communication between them is not permitted by policy.
Network infrastructure components including, routers, switches, firewalls, and other middle-boxes, are not represented explicitly as nodes. Edges represent policy-permitted end-to-end communication between end systems, abstracting away the physical or logical path taken through the network infrastructure.
Edges are assumed to be bidirectional, reflecting symmetric communication capability. 00footnotemark: 0 As $G$ is a simple graph, self-loops and parallel edges are not allowed.
Let $n=|V|$ denote the total number of nodes in the network. The maximum number of possible edges is $\binom{n}{2}$, corresponding to a fully connected network in which all pairs of nodes are allowed to communicate.
00footnotetext: We focus on segmentation within an organization’s internal network rather than perimeter security. While firewall rules can be asymmetric, our undirected edge model is appropriate for internal segmentation, as internal network security (access control) policies are often symmetric. In this paper, policy symmetry is a modeling assumption.
III-B Metric Scope
The segmentedness metric is intended to characterize policy exposure and communication permissiveness rather than architectural partitioning, network performance, or implementation details. It does not attempt to model routing or switching topology, traffic engineering, or the correctness of policy enforcement mechanisms. In addition, the metric does not capture transitive compromise or attacker capabilities beyond the existence of direct communication between two nodes, and makes no assumptions about multi-hop lateral movement or adversarial behavior beyond the ability to communicate when permitted by policy.
III-C Flatness and Segmentedness
We define flatness as a normalized measure of how permissive the network communication policy is. Specifically, the flatness of a network graph $G$ is defined as
$F(G)=\frac{|E|}{\binom{n}{2}}$, (1)
that is, the fraction of all possible node pairs for which communication is allowed.
We define segmentedness as the complement of flatness:
S(G)=1F(G).S(G)=1-F(G). (2)
Segmentedness therefore, quantifies the proportion of potential communication relationships that are disallowed by policy.
We note that the term “segmentation” is used in various ways across industry and research, with no single universally accepted definition [7]. In this paper, we use the term segmentedness to refer specifically to the metric defined above: the proportion of potential node-pair communications that are denied by policy. This operational definition allows us to provide a quantitative, comparable measure without requiring consensus on the broader conceptual boundaries of what constitutes “segmentation.”
III-D Properties
The segmentedness metric satisfies several properties useful for comparing network security policies:
  • Boundary conditions A fully connected network has segmentedness $S(G)=0$, while a fully disconnected network has segmentedness $S(G)=1$.
  • Monotonicity Adding an allowed communication edge to the network strictly decreases segmentedness, while removing an allowed edge strictly increases segmentedness.
  • Scale invariance Segmentedness is normalized and can be compared across networks of different sizes.
  • Structural neutrality Segmentedness depends only on the number of permitted communication relationships, not on their arrangement within the network.
  • Interpretability Segmentedness has a probabilistic interpretation: $S(G)$ represents the probability that two uniformly random nodes cannot communicate directly.
III-E Expected Number of Reachable Neighbors
It also turns out that for any fixed graph $G$ with edge density $F(G)$, the expected number of directly reachable neighbors of a randomly chosen node is determined solely by the edge density and the number of nodes in the network
Recall that $G=(V,E)$ is an undirected graph with $n=|V|$ nodes and that edge density
$F(G)=\frac{|E|}{\binom{n}{2}}$.
Select a node $v\in V$ uniformly at random, and let $N(v)$ denote the set of nodes that are directly reachable from $v$, i.e., the neighbors of $v$.
For any other node $u\neq v$, define an indicator random variable
$X_{u}=\begin{cases}1&\text{if }\{u,v\}\in E,\\ 0&\text{otherwise}.\end{cases}$ (3)
Then the degree of $v$ can be written as
$|N(v)|=\sum_{u\in V\setminus\{v\}}X_{u}.$ (4)
For a fixed node $v$ and a uniformly random choice of another node $u\neq v$, the probability that $\{u,v\}\in E$ equals $F(G)$, since $F(G)$ represents the fraction of all node pairs for which communication is permitted. Therefore,
$\mathbb{E}[X_{u}]=F(G)\quad\text{for all }u\neq v.$ (5)
By linearity of expectation,
$\mathbb{E}[|N(v)|]=\sum_{u\in V\setminus\{v\}}\mathbb{E}[X_{u}]=(n-1)\,F(G).$ (6)
Thus, the expected number of directly reachable neighbors of a randomly selected node is proportional to the edge density and grows linearly with the size of the network.
Since edges represent policy-permitted end-to-end communication, $\mathbb{E}[|N(v)|]=(n-1)F(G)$ can be interpreted as the expected direct exposure of a randomly chosen node: on average, a node is permitted to communicate with (n1)F(G)(n-1)F(G) other nodes. Consequently, for fixed network size $n$, reducing edge density proportionally reduces the expected number of systems that are directly reachable from any given system, which in turn reduces the expected set of immediate communication partners available to an attacker who obtains an initial foothold on a random node.
IV Empirical Estimation Methodology
We now present a practical methodology for estimating segmentedness in large networks when complete knowledge of network connectivity is unavailable. Rather than attempting to enumerate all permitted communication relationships, the approach relies on randomized sampling and empirical connectivity testing to estimate segmentedness with statistical confidence.
IV-A Measurement Challenge
Direct computation of segmentedness requires knowledge of all permitted communication pairs in the network, which is infeasible in large enterprise environments. For a network with $n$ nodes, there are $\binom{n}{2}$ possible node pairs, making exhaustive testing impractical even for moderately sized networks. In addition, network security policies are often distributed across systems, further complicating direct analysis.
As a result, segmentedness must be estimated empirically using partial observations, while providing explicit bounds on estimation error.
IV-B Randomized Sampling of Node Pairs
To estimate $F(G)$, we randomly sample $M$ unordered pairs of distinct nodes u,v{u,v} from $V$ with replacement. For each sampled pair, we perform a connectivity test to determine whether communication between $u$ and $v$ is permitted by policy. Each test yields a Bernoulli outcome:
$X_{i}=\begin{cases}1&\text{if communication is permitted},\\ 0&\text{otherwise}.\end{cases}$ (7)
The empirical flatness estimate is then given by
$\hat{F}=\frac{1}{M}\sum_{i=1}^{M}X_{i},$ (8)
with corresponding segmentedness estimate $\hat{S}=1-\hat{F}$.
IV-C Connectivity Test Suite
Determining whether communication between two nodes is permitted requires operational testing. Relying on a single test, such as ICMP echo, can produce false negatives due to protocol-specific filtering or configuration. To improve robustness, we propose a battery of connectivity tests that may include ICMP probes, UDP packet exchanges, and TCP connection attempts.
A node pair is considered connected if any test in the suite succeeds. This definition reflects the security-relevant observation that the existence of at least one usable communication channel is sufficient for interaction between systems. The test suite is fixed for all measurements to ensure consistency and comparability across sampling runs.
IV-D Statistical Guarantees and Sample Size
Each sampled pair test can be viewed as an independent Bernoulli trial with success probability $p=F(G)$. Under this model, the estimator $\hat{F}$ is an unbiased estimator of $p$ with variance
$\mathrm{Var}(\hat{F})=\frac{p(1-p)}{M}.$ (9)
By the central limit theorem [8], for sufficiently large $M$ (typically $M\geq 30$ suffices), $\hat{F}$ is approximately normally distributed. An approximate two-sided confidence interval with confidence level $1-\alpha$ is given by
$\hat{F}\pm z_{\alpha/2}\sqrt{\frac{p(1-p)}{M}},$ (10)
where $z_{\alpha/2}$ is the corresponding quantile of the standard normal distribution. This is known as the Wald or normal approximation confidence interval [1]. 00footnotemark: 0
00footnotetext: We use the Wald confidence interval because it yields a simple, closed-form sample size requirement and enables conservative, network-size-independent guarantees under worst-case variance assumptions.
Since $p$ is unknown, a conservative bound can be obtained by noting that p(1p)p(1-p) is maximized at $p=0.5$. This yields a worst-case standard error of
$\mathrm{SE}(\hat{F})\leq\frac{1}{2\sqrt{M}}.$ (11)
This worst-case assumption means that the required sample size $M$ is valid for any network, regardless of its actual segmentedness level.
To achieve an absolute error of at most $\varepsilon$ with confidence level $1-\alpha$, it suffices to choose
$M\geq\frac{z_{\alpha/2}^{2}}{4\varepsilon^{2}}.$ (12)
Because segmentedness is defined as S(G)=1F(G)S(G)=1-F(G), the same confidence interval and error bounds apply directly to $\hat{S}$. Importantly, the required sample size depends only on the desired accuracy and confidence level, and not on the total number of nodes in the network, provided that node pairs are sampled uniformly at random.
V Example Sample Size Calculations
Using the formula derived in Section IV-D, we compute the minimum number of randomly sampled node pairs $M$ required to estimate segmentedness with a specified confidence interval. We consider a two-sided 95% confidence interval (CI) with half-width $\varepsilon=0.1$.
For a 95% confidence level, the corresponding normal quantile is $z_{0.025}\approx 1.96$. Applying the conservative bound that maximizes variance at $p=0.5$, the required sample size satisfies
$M\geq\frac{z_{0.025}^{2}}{4\varepsilon^{2}}.$ (13)
Substituting the desired values yields
$M\geq\frac{(1.96)^{2}}{4(0.1)^{2}}=\frac{3.8416}{0.04}=96.04.$ (14)
Rounding up to the nearest integer, this gives a minimum of $M=97$ sampled node pairs. Notably, this result is independent of the total number of nodes in the network, provided that node pairs are sampled uniformly at random.
Table I summarizes the corresponding conservative estimates of $M$ for networks with 10,000 and 100,000 nodes under the same confidence requirements. Tables II and III show the estimates for $M$ for networks with 1,000, 10,000 and 100,000 nodes, when the confidence coefficient and the half-width are varied.
That a 95% CI with a half-width of $\pm 0.1$ can be achieved with just $97$ node pairs is a nice result—it makes the use of edge density as a measure of segmentedness feasible for networks of practically arbitrary size. Segmentedness measurements can be repeated frequently (for example, every quarter), even if they have to be run manually.
TABLE I: Conservative minimum number of randomly sampled node pairs $M$ required to estimate segmentedness with a 95% CI of $\pm 0.1$, shown alongside the maximum number of possible edges for each network size.
# nodes Max. possible edges Min. $M$ 95% CI $\pm 0.1$
1,0001{,}000 499,500499{,}500 $97$
10,00010{,}000 49,995,00049{,}995{,}000 $97$
100,000100{,}000 4,999,950,0004{,}999{,}950{,}000 $97$
TABLE II: Conservative minimum number of randomly sampled node pairs $M$ required to estimate segmentedness with a CI of $\pm 0.1$ at 90%, 95%, and 99% confidence levels.
# nodes $M$ (90% CI) $M$ (95% CI) $M$ (99% CI)
1,0001{,}000 $68$ $97$ $166$
10,00010{,}000 $68$ $97$ $166$
100,000100{,}000 $68$ $97$ $166$
TABLE III: Conservative minimum number of randomly sampled node pairs $M$ required to estimate segmentedness with a CI of $\pm 0.05$ at 90%, 95%, and 99% confidence levels.
# nodes $M$ (90% CI) $M$ (95% CI) $M$ (99% CI)
1,0001{,}000 $271$ $385$ $664$
10,00010{,}000 $271$ $385$ $664$
100,000100{,}000 $271$ $385$ $664$
VI Evaluation via Simulations
We evaluate the proposed estimation procedure through Monte Carlo simulations on synthetic networks. We begin with the Erdős–Rényi model to establish baseline estimator behavior, then assess robustness under the stochastic block model, which introduces heterogeneous connectivity patterns representative of segmented enterprise networks.
VI-A Estimation over Random Networks
We begin by evaluating the proposed estimation procedure in a simple and well-understood random network setting. The Erdős–Rényi model provides a natural baseline, as it captures networks in which connections between nodes occur independently and uniformly at random, while still allowing the overall level of connectivity to be controlled by a single parameter. This makes it a useful reference point for understanding estimator behavior before considering more structured graph models.
In the Erdős–Rényi model $G(n,p)$ [4], a graph is generated on $n$ labeled nodes by including each unordered pair of distinct nodes as an edge independently with probability $p$. In this setting, the edge probability coincides with the edge density $F(G)$. Rather than explicitly constructing the full graph, the estimation procedure samples $M$ unordered node pairs uniformly at random and observes whether an edge is present. Each observation corresponds to a Bernoulli random variable $X_{i}\sim\mathrm{Bernoulli}(F(G))$, and the estimator $\hat{F}=\frac{1}{M}\sum_{i=1}^{M}X_{i}$ coincides with the empirical edge density defined earlier.
Figure 1 illustrates the Monte Carlo behavior of the estimator $\hat{F}$ for Erdős–Rényi graphs with n=100,000n=100{,}000 nodes, chosen to mimic the scale of a large organizational network. The true edge probability $p$ is varied over the range $0.1\leq p\leq 0.5$, corresponding to different levels of network connectivity. For each value of $p$, the markers show the mean estimate of $\hat{F}$ across repeated (1,0001{,}000) trials, and the shaded region indicates the associated $95\%$ confidence interval. The estimates lie close to the diagonal across all values of $p$, confirming that the estimator is unbiased in this setting and behaves stably.
Figure 2 reports the empirical coverage probability of the $95\%$ Wald confidence interval for the same Erdős–Rényi simulations. For each value of $p$, coverage is computed as the fraction of Monte Carlo trials in which the confidence interval contains the true edge density. Across the full range of connectivity levels considered, the observed coverage remains close to the nominal level of $0.95$, indicating that the confidence intervals used in this paper are well-behaved under the Erdős–Rényi baseline.
Figure 1: Monte Carlo mean and 95% CI of the edge density estimator under the Erdős–Rényi $G(n,p)$ model as a function of the true edge probability $p$.
Figure 2: Empirical coverage probability of the 95% Wald confidence interval for the edge density estimator under the Erdős–Rényi $G(n,p)$ model as a function of the true edge probability $p$.
VI-B Estimation over Networks with Equal-sized Communities
We next evaluate the proposed estimation procedure in a random network model with explicit community structure. The stochastic block model (SBM) provides a natural extension of the Erdős–Rényi baseline by allowing edge probabilities to vary across different classes of node pairs [10]. This setting makes it possible to assess the robustness of the estimator when global edge density is controlled by local connectivity patterns.
In the SBM with $K$ equal-sized blocks, a graph is generated on $n$ labeled nodes by first partitioning the nodes into $K$ disjoint groups of equal size. Edges between nodes in the same block are included independently with probability $p_{\text{in}}$, while edges between nodes in different blocks are included independently with probability $p_{\text{out}}$. Although all edge decisions remain independent, the probability of an edge now depends on the block memberships of the node pair. The resulting expected edge density is a weighted average of $p_{\text{in}}$ and $p_{\text{out}}$:
$F(G)=\frac{1}{K}p_{\text{in}}+\frac{K-1}{K}p_{\text{out}},$ (15)
where the weights $\frac{1}{K}$ and $\frac{K-1}{K}$ represent the proportions of within-block and between-block node pairs, respectively.
As in the Erdős–Rényi setting, the estimation procedure does not require construction of the full graph. Instead, $M$ unordered node pairs are sampled uniformly at random, and an indicator variable is recorded for the presence or absence of an edge. Each sampled indicator follows a Bernoulli distribution with success probability equal to the global edge density of the SBM, and the estimator $\hat{F}$ is defined as the sample mean of these observations.
Figures 3 and 4 illustrate the Monte Carlo behavior of the estimator $\hat{F}$ for SBM networks with n=100,000n=100{,}000 nodes and $K=5$ equal-sized blocks. In each figure, the between-block probability $p_{\mathrm{out}}$ is held fixed, with $p_{\mathrm{out}}=0.1$ in Figure 3 and $p_{\mathrm{out}}=0.2$ in Figure 4, while the within-block probability $p_{\mathrm{in}}$ is varied to produce global edge densities in the range $[0.1,0.5]$, similar to Section VI-A. For each parameter configuration, the markers show the mean of $\hat{F}$ across 1,0001{,}000 Monte Carlo trials, and the shaded region represents the associated $95\%$ confidence interval. Across the full range of edge densities considered, the estimates closely follow the identity line, indicating that the estimator remains approximately unbiased in the presence of community structure.
Figure 5 reports the empirical coverage probability of the $95\%$ Wald confidence interval for the same SBM simulations. Coverage is computed as the fraction of Monte Carlo trials in which the confidence interval contains the true global edge density. For both values of $p_{\mathrm{out}}$, the observed coverage remains close to the nominal level of $0.95$ across the range of edge densities examined. These results suggest that the Wald confidence intervals used in this paper remain well-behaved for edge density estimation under equal-sized stochastic block models.
The SBM results demonstrate that the sampling-based estimation procedure remains accurate even when network structure exhibits significant heterogeneity, as would be expected in real enterprise networks with business-unit or functional segmentation. Although the segmentedness metric itself is structurally neutral and depends only on global edge density (Section III-D), these results confirm that the estimation procedure works correctly regardless of how edges are arranged within the network. Importantly, because the estimation procedure samples node pairs uniformly at random without constructing the full graph, the estimator’s behavior depends only on the global edge density and is invariant to the underlying network topology. Validation on alternative graph models, such as small-world or scale-free networks, produces identical results for the same edge density, confirming the structural independence of the approach.
Figure 3: Monte Carlo mean and 95% CI of the edge density estimator under the SBM equal-sized blocks model (p_out=0.1) as a function of the true global edge density $p$.
Figure 4: Monte Carlo mean and 95% CI of the edge density estimator under the SBM equal-sized blocks model (p_out=0.2) as a function of the true global edge density $p$.
Figure 5: Empirical coverage probability of the 95% Wald confidence interval for the edge density estimator under the SBM equal-sized model as a function of the true global edge density $p$.
VII Operational Use Cases
With a well-defined metric in hand, several use cases are possible.
Baseline measurement and trend tracking A primary operational use case is establishing a baseline segmentedness measurement and tracking changes over time. For example, an organization with 10,00010{,}000 endpoints might conduct quarterly measurements using $M=97$ sampled pairs ($95\%$ CI, $\pm 0.1$ half-width). If the initial measurement yields $\hat{S}=0.3(\pm 0.1)$, indicating $30\%$ of potential communications are blocked, subsequent measurements can track whether segmentation initiatives are improving network isolation. A decrease to $\hat{S}=0.2$ in the following quarter would indicate policy drift or new connectivity requirements reducing segmentation levels, prompting investigation.
Environment consolidation Organizations frequently consolidate previously separate environments, such as development, staging, and production networks, or merge multiple data centers into a shared infrastructure platform. These initiatives often introduce broad new communication paths between systems that were previously isolated, resulting in large changes to network policy permissiveness. Segmentedness measurements taken before and after consolidation can quantify the security impact of these architectural decisions and help ensure that consolidation efforts do not unintentionally produce an overly flat network.
Merger and acquisition integration During mergers or acquisitions, organizations must integrate networks while maintaining security boundaries. The segmentedness metric provides a quantitative measure of integration progress. For instance, two networks with pre-merger segmentedness values of $\hat{S_{1}}=0.5$ and $\hat{S_{2}}=0.4$ might temporarily show $\hat{S}_{combined}=0.3$ immediately after connection, reflecting increased connectivity between previously isolated systems. Tracking this metric helps security teams ensure that integration does not inadvertently create an overly flat combined network.
[…]
Benchmarking While no established benchmarks currently exist for segmentedness values, the metric enables future development of industry-specific or sector-specific guidelines. Financial services organizations, which typically enforce stricter isolation requirements, might target $\hat{S}\geq 0.6$, while educational institutions with more open collaboration requirements might operate effectively at $\hat{S}\approx 0.3$. As organizations adopt this metric, empirical data will enable peer comparison and the establishment of reasonable target ranges.
VIII Practical Considerations for Sampling Node Pairs
The segmentedness metric assumes that node pairs are sampled uniformly at random from the set of networked entities. In operational environments, implementing such sampling introduces practical challenges that may affect how closely real-world measurements adhere to this assumption. While Section IX discusses sampling limitations from a statistical perspective, this section focuses on operational challenges that practitioners may encounter when deploying the measurement methodology.
Incomplete node enumeration Uniform random sampling presumes the ability to enumerate the set of nodes over which communication policies are defined. In practice, asset inventories may be incomplete or dynamic, particularly in environments with ephemeral virtual machines, containers, or unmanaged endpoints. As a result, sampling procedures may disproportionately select long-lived or well-managed systems, potentially biasing measurements toward portions of the network with more explicit or carefully maintained policies.
Tooling-induced sampling bias In many organizations, sampling node pairs is mediated by security or management tooling, such as firewall policy managers and configuration databases. These tools often expose subsets of the network more readily than others, for example, privileging server-to-server communication over user endpoints or unmanaged devices. Consequently, node pairs may be sampled from the set of entities visible to the tooling rather than uniformly from the underlying network, introducing visibility-driven bias.
Configured vs enforced policy The metric is defined in terms of configured policies, but measurements in practice reflect observed enforcement. The two may differ due to routing constraints, implicit deny rules, or configuration drift.
Organizational and administrative constraints Beyond technical considerations, organizational factors may restrict which node pairs can be evaluated. Access controls, change management processes, or business unit boundaries may limit the ability to probe certain communication paths. As a result, practical sampling procedures may be partially randomized and partially constrained by administrative policy.
Taken together, these considerations highlight that sampling node pairs in real networks is an engineering challenge rather than a purely statistical one. The segmentedness metric remains informative under such constraints when tracking relative changes over time, particularly when large-scale policy modifications (e.g., zero trust deployment, merger integration) produce measurable shifts in overall permissiveness that exceed the bias introduced by sampling constraints.
IX Limitations
The proposed metric and methodology have several important limitations that should be considered when interpreting results.
Binary connectivity model The metric treats communication as binary (permitted/denied) without capturing degrees of restriction. For example, connections limited to specific ports or protocols are treated identically to unrestricted connections. Future work could incorporate connection constraints into a weighted edge model.
Static measurement The metric captures policy-permitted connectivity at a point in time and does not reflect dynamic access control mechanisms, temporal restrictions, or context-dependent policies. Networks using adaptive security models may require repeated measurements under different conditions.
Symmetric edge assumption As noted in Section III-A, the undirected edge model assumes policy symmetry. Networks with significant asymmetric policies (e.g., client-server architectures with one-way initiation requirements) may not be fully represented.
Topology independence While structural neutrality is a design goal (Section 3.4), it means the metric does not capture important security-relevant properties such as hub-and-spoke vulnerabilities, critical chokepoints, or segmentation hierarchy.
Sampling limitations The statistical guarantees assume uniform random sampling and independent connectivity tests. Non-uniform sampling strategies—such as oversampling from specific subnets or sampling geographically clustered nodes—could introduce correlation between samples and bias estimates. Organizations should ensure sampling is truly random across the entire endpoint population. See Section VIII for a detailed discussion of practical sampling challenges in operational environments.
Perimeter vs. internal The metric focuses on internal network segmentation and does not characterize perimeter security, Internet connectivity, or external threat exposure.
X Conclusions
We introduce segmentedness as a simple scalar metric for network isolation. By formalizing segmentedness through edge density, we provide practitioners with an interpretable measure of how permissive network communication policies are.
The metric satisfies key properties for practical use: it is normalized for cross-network comparison, scales to networks of arbitrary size, and can be estimated efficiently through randomized sampling with explicit statistical guarantees. Our validation demonstrates that estimates remain unbiased and well-behaved across different network structures.
The main contribution of this paper is in providing a quantitative measure that can be compared across networks and over time. Organizations can now track segmentation trends (e.g., through quarterly measurements requiring fewer than 100 samples), validate security architecture initiatives, and establish empirical baselines for their networks.
Future work includes (a) empirical studies across diverse organizational networks to establish sector-specific benchmarks (e.g., financial services, healthcare, education), (b) extensions incorporating connection constraints and policy attributes beyond binary connectivity, (c) integration with automated network management and security management platforms, and (d) development of complementary structural metrics to provide a comprehensive segmentation assessment.
As organizations continue adopting zero trust architectures and micro-segmentation strategies, quantitative metrics like segmentedness will become essential tools for measuring progress, demonstrating compliance, and making data-driven security decisions.